six question multiple choice statistics exam is given. Each question has 4 possible choices, only one if which is correct. Therefore, the probability of success for each exam is 1/4 or 0.25. What is the probability of passing the exam (4 or more questions), by randomly guessing the answers?
Answer by Boreal(15235)

It is the probability of answering 4 questions, 5 questions or 6 questions correctly.
probability of 4 is 6C4(0.75)^2(0.25)^4
probability of 5 is +(0.75)(0.25)^5
probability of 6 is 0.25^6=
0.0330+0.004+0.0002=0.0334